# aboveskip=2pt, belowskip=2pt

### v1

@@ Abstract
The Arrow of Time (AoT)—time’s irreversible flow shaping physical events—is fundamental to video comprehension, yet remains a significant challenge for modern large multimodal models (LMMs). Current LMMs struggle to perceive and utilize temporal directionality in video when responding to language queries, obstructing deeper temporal understanding. We tackle this deficiency by first providing a critical analysis of existing benchmarks and models. We then introduce ArrowRL, a reinforcement learning (RL)-based training strategy with an innovative reverse reward that instills AoT awareness by encouraging divergent video interpretations between forward and reversed visual frames. For rigorous evaluation, we additionally develop AoTBench, a new multi-faceted benchmark probing temporally challenging questions. Experiments show ArrowRL greatly advances temporal perception: it not only achieves substantial improvements on our challenging AoTBench but also demonstrably boosts performance on standard video question answering (VQA) benchmarks (with peak accuracy gains reaching over 20% and 10% respectively). This validates ArrowRL’s effectiveness and highlights the critical need for dedicated AoT understanding in LMMs. [^1] Project webpage: https://vision.cs.utexas.edu/projects/SeeAoT.
@@ 1 Introduction
Naturally, this leads us to ask: how capable are current LMMs in perceiving the AoT? Perhaps surprisingly, despite rapid advancements elsewhere, our findings indicate a critical gap. First, a profound temporal insensitivity is evident even in state-of-the-art LMMs. They fail on basic AoT-related tasks, such as distinguishing forward from backward video playback (Fig. 1 (a)), and ignore crucial semantic differences between forward and reverse sequences, erroneously providing identical descriptions (Fig. 1 (b)). Second, broadening our investigation, we identify an alarming failure: on multiple standard video question-answering (VQA) benchmarks xiao2021nextqa ; fu2024videomme ; patraucean2023perception ; mangalam2023egoschema ; cai2024temporalbench ; li2024vitatecs ; li2024mvbench ; wu2024longvideobench , shuffling or reversing video frames often causes only slight or even no performance degradation for a leading representative LMM (see Fig. 2 and details in Supp.).
@@ 2 Related Work
Time has long been recognized as a valuable self-supervised signal in video pickup2014seeingaot ; wei2018learningaot ; ghodrati2018videotime ; misra2016shuffle ; dorkenwald2022scvrl ; lee2017unsupervised ; jing2018self ; dwibedi2019temporal ; xue2023learning . Early works capitalize on this property within the visual domain, leveraging the arrow of time (AoT) by using forward/reverse sequence classification as a pretraining task, which benefits downstream action recognition pickup2014seeingaot ; wei2018learningaot ; ghodrati2018videotime . Relatedly, random shuffling has been explored for action analysis misra2016shuffle ; price2019retro ; sevilla2021only . Current multimodal vision foundation models gan2022vision , such as CLIP radford2021learning and VideoCLIP xu2021videoclip , now benefit action understanding, driven by text-video contrastive alignment objectives. A few works bagad2023test ; wang2023paxion ; du2024reversed employ temporal reversal as a source of negative examples for aligning video and text encoders, yet the discussion remains specific to dual-encoder architectures and only addresses basic temporal concepts with before/after relations bagad2023test or action antonyms wang2023paxion ; du2024reversed . Critically, while the field has advanced to LMMs capable of free-form language-based interaction with video content liang2024survey ; caffagni2024revolution , we find that these models exhibit a surprising insensitivity to fundamental temporal directionality. This suggests their internal world models hao2023reasoning ; ha2018world ; mendonca2023structured may fail to incorporate the basics of real-world event progression and causality. How to instill AoT awareness within LMMs remains a key unexplored area.
@@
LLM capabilities have been extended to the visual domain through the development of LMMs liang2024survey ; caffagni2024revolution , which integrate a vision encoder with a powerful LLM, enabling them to perceive and reason about visual content. While early LMMs focus on images liu2023visual ; liu2024improved ; zhu2023minigpt ; tong2024cambrian , research on video-LMMs chen2024sharegpt4video ; zhang2024llavavideo ; zohar2024apollo ; li2025exploring ; li2024llavaov ; wang2024qwen2 ; bai2025qwen25 ; wang2024tarsier ; liu2024st ; xu2024pllava ; cho2025perceptionlm is rapidly progressing, including the development of high-quality video datasets chen2024sharegpt4video ; zhang2024llavavideo and exploration of the architecture design space zohar2024apollo ; li2025exploring ; liu2024st . Our work is orthogonal to these data and architecture innovations. We investigate how a core understanding of the AoT can be instilled directly through training, offering a complementary path towards enhanced temporal perception in LMMs.
@@
Post-training using reinforcement learning (RL) is increasingly employed to refine LLMs, aligning model responses with human preferences and enhancing complex reasoning abilities ouyang2022training ; ziegler2019fine ; rafailov2023direct ; guo2024direct ; shao2024deepseekmath . A family of successful algorithms, including PPO ziegler2019fine , DPO rafailov2023direct , online DPO guo2024direct , and GRPO shao2024deepseekmath , is now being actively extended to LMMs, although applications have mainly focused on image-LMMs sun2023aligning ; yu2024rlhf ; zhang2025r1 ; zhang2024improve , with video comprehension being a newer frontier. While these efforts have concentrated on mitigating visual hallucination zhang2024direct ; wang2024videohallucer ; zhang2024eventhallusion or (concurrently) enabling structured reasoning feng2025videor1 ; chen2025exploring like Chain-of-Thought wei2022chain , our work investigates a distinct dimension: enhancing the model’s fundamental temporal awareness, by leveraging AoT signals naturally embedded in videos.
@@ 3.1 Are Today’s Benchmarks Time-Sensitive? Revisiting Current LMM Evaluation
We then evaluate the three selected LMMs across eight popular VQA benchmarks (VITATECS li2024vitatecs , TemporalBench cai2024temporalbench , NExT-QA xiao2021nextqa , PerceptionTest patraucean2023perception , VideoMME fu2024videomme , Vinoground zhang2024vinoground , TempCompass liu2024tempcompass and TVBench cores2024tvbench ). [^2] Note that our analysis focuses on MCQ tasks for their popularity and clear evaluation structure, as opposed to open-ended QA or captioning tasks that require a third-party LLM evaluator, introducing additional uncertainty. To quantify benchmark and sample-level AoT sensitivity (facilitating later construction of AoTBench, c.f., Sec. 3.3), we propose the temporal divergence score (TDS). For an evaluator LMM $f$, consider an MCQ sample $i$, comprised of video $v_{i}$, language query $l_{i}$ and ground truth answer $y_{i}$ from $K_{i}$ options. Denote the reversed video as $\tilde{v}_{i}$. We compute the first output token probability distributions over the $K_{i}$ options, for both forward and reversed video: $p_{i}=\mathrm{TokenProb}(f,v_{i},l_{i})$ and $\tilde{p}_{i}=\mathrm{TokenProb}(f,\tilde{v}_{i},l_{i})$, where $p_{i},\tilde{p}_{i}\in\mathbb{R}^{K_{i}}$. We first filter out instances where the average reverse performance exceeds its forward performance. The TDS for any remaining sample $i$ is defined using KL divergence: $\mathrm{TDS}_{i}=\mathbb{D}_{\text{KL}}[p_{i}\|\tilde{p}_{i}]$. A high TDS signifies that the model’s prediction is highly sensitive to the video’s temporal direction—causing disagreement between forward and reverse responses—therefore suggesting the sample requires AoT understanding for a correct interpretation and answer.
@@ 3.2 ArrowRL: Learning the Arrow of Time
We now discuss our approach to equipping a given LMM $\pi_{\theta}$ with AoT understanding for enhanced temporal perception. Our core premise is that videos are inherently dynamic, and LMMs should capture this dynamic progression when responding. Specifically, given a question $q=(v,l)$ (composed of video $v$ and language query $l$), the goal is for $\pi_{\theta}$ to produce responses that align well with a target response $o^{\star}$, operating on standard instruction tuning datasets $P(Q,O^{\star})$. To foster broad AoT sensitivity, our training utilizes a diverse suite of tasks—MCQ, open-ended QA, and video captioning (see Sec. 4 for the specific datasets used).
@@
First, we define a fidelity reward encouraging alignment between each candidate response $o_{i}$ and the target response $o^{\star}$, $r_{i}^{\mathrm{\ fid}}=\mathrm{Similarity}(o_{i},o^{\star})$. The function $\mathrm{Similarity}(\cdot,\cdot)$ returns a scalar between 0 and 1 and is calculated via accuracy matching (1.0 if $o_{i}=o^{\star}$, else 0.0) for MCQ tasks, or via an LLM-based similarity score for open-ended QA and captioning (see Supp. for the specific prompts and the LLM judge employed).
@@
Second, central to ArrowRL is the reverse reward. Rewarding only fidelity to the target response $o^{\star}$ (via $r_{i}^{\mathrm{\ fid}}$) can be insufficient for capturing directional nuance—particularly if $o^{\star}$ lacks temporal progression details. Therefore, we utilize the reversed video (denoted by $\tilde{v}$) as a natural directional contrast signal, which often distinguishes itself from the forward version, to enforce AoT sensitivity. Intuitively, an AoT-sensitive model should yield different responses when observing the same video in different sequential order (though not without exceptions, which we handle below). Specifically, we maximize the dissimilarity between forward candidate responses $o_{i}$ and the reverse response $\tilde{o}$, i.e., the response produced by $\pi_{\theta}$ when given $\tilde{v}$. The reverse reward is thus defined as: $r_{i}^{\mathrm{\ rev}}=1-\mathrm{Similarity}(o_{i},\tilde{o})$. This reward design discourages generic or static responses and compels attention to the dynamic, directional details conveyed by AoT in forward $v$.
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\mathbb{E}_{(q,o^{\star})\sim P(Q,O^{\star}),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}\hat{A}_{i,t},$
@@ 4 Experiments
Our ArrowRL training data comprises a comprehensive suite of tasks, with data selected or adapted for each to emphasize scenarios requiring AoT awareness: (1) MCQ tasks: we format sequence direction classification (Fig. 1 (a)) as MCQ, using 1.1K training videos selected from UCF101 soomro2012ucf101 (following prior use of this dataset ghodrati2018videotime ); (2) Open-ended QA: we curate a high-temporality subset of LLaVA-Video-178K zhang2024llavavideo , filtering based on the perplexity difference between forward and reverse video to retain samples with great temporal sensitivity, totaling 11.8K samples; (3) Video Captioning: we employ the training set du2024reversed of RTime, which provides high-temporality videos alongside distinct human captions for their forward and reverse versions, comprising 11.7K samples. See Supp. for data examples and prompts used.
@@
Our framework presents a straightforward and highly efficient approach by enabling direct application to pretrained LMMs without an intermediate supervised fine-tuning (SFT) step; we find ArrowRL alone is sufficient to instill AoT awareness. We demonstrate its broad applicability across three leading base LMMs: LLaVA-OV-7B li2024llavaov , Qwen2-VL-7B wang2024qwen2 , and Qwen2.5-VL-7B bai2025qwen25 . For efficiency and memory, we limit the maximum number of video frames to 16 during training. The hyperparameters balancing the reward ($\alpha$) and controlling the dynamic weighting ($\gamma$) are both set to 0.5. The response group size $G$ is set as 8. Training consists of 2000 RL steps on 6 NVIDIA GH200 GPUs.
@@
Results
As shown in Table 1, ArrowRL effectively enhances models’ temporal perception capabilities across diverse settings. Results on AoTBench (left table) reveal critical deficiencies in current LMM AoT perception. Visually-focused tasks like sequence direction classification prove especially challenging: nearly all open-source models exhibit a universal failure mode: their responses are identical regardless of video input, resulting in consistent chance-level performance (50.0% on ReverseFilm and 51.6% on UCF101). Leading proprietary models (GPT-4o, Gemini-1.5-Pro), while slightly better, still perform far below satisfactory levels. This is in stark contrast to the proprietary models’ absolute accuracy on other mainstream VQA tasks, where accuracy hovers around 70-85%. Our finding and the proposed AoTBench illuminate visual temporal bottlenecks. Importantly, this improved AoT awareness also translates to gains on existing, unfiltered VQA benchmarks (right table), such as +10.8 group score (+65.9% relative gain) for Qwen2.5-VL-7B on Vinoground. These results underscore the promise of enhancing AoT perception for overall temporal perception. See Supp. for a follow-up discussion of the relative strengths of different base models and how they translate to ArrowRL’s largest margins.
@@ Results
Table 2: Ablation results of ArrowRL using Qwen2.5-VL-7B as the base LMM. We report average performance across all columns of AoTBench tasks in Table 1 (see Supp. for detailed per-task results).
Model Acc. (%)
Qwen2.5-VL-7B bai2025qwen25 56.2 $\downarrow$
[…]
+ ArrowRL ($\gamma=1$) 58.3 $\uparrow$
+ SFT 57.4 $\uparrow$
[…]
Further Analysis of ArrowRL
Table 2 presents the ablation study of ArrowRL. We observe the critical role of the reverse reward: removing it (i.e., setting its default weight $\alpha=0$) degrades the performance below base. Furthermore, disabling our dynamic weighting scheme (i.e., setting the threshold $\gamma=1$) yields less improvement than allowing $\alpha_{i}$ to dynamically become zero for non-AoT-sensitive samples. Finally, we experiment with standard SFT on the same training data. While SFT improves over the base model (suggesting our training data quality), ArrowRL achieves larger gains, underscoring the effectiveness of our RL approach. We further investigate different inference frame settings and find ArrowRL generalizes robustly despite its fixed-frame training. See Supp. for details.
Figure 7: Qualitative results comparing base LMMs vs. Arrow-RL enhanced models. ArrowRL correctly answers AoT-sensitive VQA questions (left) and generates distinct, temporally accurate captions (right), while the base model fails to capture temporal progression in these examples. More examples and failure cases in Supp.
@@
Fig. 7 provides qualitative examples, comparing base LMMs against our ArrowRL enhancement. The VQA examples (left) demonstrate ArrowRL enabling base models to correctly answer temporally challenging questions. Similarly, when prompted to describe forward and reverse sequences (right), the base model provides “unwrapping” description for both, failing to connect frames dynamically to understand the temporal progression and possibly influenced by language bias (favoring the more common “unwrapping”). Conversely, ArrowRL demonstrates clear temporal awareness by producing distinct and correct captions reflecting the temporal direction. See Supp. for more qualitatives including failure cases and a comprehensive discussion of limitations.

### v2

# Seeing the Arrow of Time in Large Multimodal Models

@@ Abstract
The Arrow of Time (AoT)—time’s irreversible flow shaping physical events—is fundamental to video comprehension, yet remains a significant challenge for modern large multimodal models (LMMs). Current LMMs struggle to perceive and utilize temporal directionality in video when responding to language queries, obstructing deeper temporal understanding. We tackle this deficiency by first providing a critical analysis of existing benchmarks and models. We then introduce ArrowRL, a reinforcement learning (RL)-based training strategy with an innovative reverse reward that instills AoT awareness by encouraging divergent video interpretations between forward and reversed visual frames. For rigorous evaluation, we additionally develop AoTBench, a new multi-faceted benchmark probing temporally challenging questions. Experiments show ArrowRL greatly advances temporal perception: it not only achieves substantial improvements on our challenging AoTBench but also boosts performance on standard video question answering (VQA) benchmarks (with peak accuracy gains reaching over 20% and 10% respectively). This validates ArrowRL’s effectiveness and highlights the critical need for dedicated AoT understanding in LMMs. [^1] Project webpage: https://vision.cs.utexas.edu/projects/SeeAoT.
@@ 1 Introduction
Naturally, this leads us to ask: how capable are current LMMs in perceiving the AoT? Perhaps surprisingly, despite rapid advancements elsewhere, our findings indicate a critical gap. First, a profound temporal insensitivity is evident even in state-of-the-art LMMs. They fail on basic AoT-related tasks, such as distinguishing forward from backward video playback (Fig. 1 (a)), and ignore crucial semantic differences between forward and reverse sequences, erroneously providing identical descriptions (Fig. 1 (b)). [^2] This is not a general failure to process reversed videos as out-of-distribution (OOD) inputs. See Supp. B.1 for a control experiment where LMMs excel on a temporally-insensitive task regardless of video direction. Second, broadening our investigation, we identify an alarming failure: on multiple standard video question-answering (VQA) benchmarks xiao2021nextqa ; fu2024videomme ; patraucean2023perception ; mangalam2023egoschema ; cai2024temporalbench ; li2024vitatecs ; li2024mvbench ; wu2024longvideobench , shuffling or reversing video frames often causes only slight or even no performance degradation for a leading representative LMM (see Fig. 2 and details in Supp. B.2).
@@ 2 Related Work
Time has long been recognized as a valuable self-supervised signal in video pickup2014seeingaot ; wei2018learningaot ; ghodrati2018videotime ; misra2016shuffle ; dorkenwald2022scvrl ; lee2017unsupervised ; jing2018self ; dwibedi2019temporal ; xue2023learning . Early works capitalize on this property within the visual domain, leveraging the arrow of time (AoT) by using forward/reverse sequence classification as a pretraining task, which benefits downstream action recognition pickup2014seeingaot ; wei2018learningaot ; ghodrati2018videotime . Relatedly, random shuffling has been explored for action analysis misra2016shuffle ; price2019retro ; sevilla2021only . Current multimodal vision foundation models gan2022vision , such as CLIP radford2021learning and VideoCLIP xu2021videoclip , can benefit action understanding, driven by text-video contrastive alignment objectives. A few works bagad2023test ; wang2023paxion ; du2024reversed employ temporal reversal as a source of negative examples for aligning video and text encoders, yet the discussion remains specific to dual-encoder architectures and only addresses basic temporal concepts with before/after relations bagad2023test or action antonyms wang2023paxion ; du2024reversed . Critically, while the field has advanced to LMMs capable of free-form language-based interaction with video content liang2024survey ; caffagni2024revolution , we find that these models exhibit a surprising insensitivity to fundamental temporal directionality. This suggests their internal world models hao2023reasoning ; ha2018world ; mendonca2023structured may fail to incorporate the basics of real-world event progression and causality. While recent work investigates AoT in the context of text-only LLMs papadopoulos2024arrows , and concurrent efforts explore it for physical reasoning azzolini2025cosmos , how to instill AoT awareness within LMMs remains a critical but largely unexplored problem.
@@
LLM capabilities have been extended to the visual domain through the development of LMMs liang2024survey ; caffagni2024revolution , which integrate a vision encoder with a powerful LLM, enabling them to perceive and reason about visual content. While early LMMs focus on images liu2023visual ; liu2024improved ; zhu2023minigpt ; tong2024cambrian , research on video-LMMs chen2024sharegpt4video ; zhang2024llavavideo ; zohar2024apollo ; li2025exploring ; li2024llavaov ; wang2024qwen2 ; bai2025qwen25 ; wang2024tarsier ; liu2024st ; xu2024pllava ; cho2025perceptionlm ; xue2025progress is rapidly progressing, including the development of high-quality video datasets chen2024sharegpt4video ; zhang2024llavavideo and exploration of the architecture design space zohar2024apollo ; li2025exploring ; liu2024st . Our work is orthogonal to these data and architecture innovations. We investigate how a core understanding of the AoT can be instilled directly through training, offering a complementary path towards enhanced temporal perception in LMMs.
@@
Post-training using reinforcement learning (RL) is increasingly employed to refine LLMs, aligning model responses with human preferences and enhancing complex reasoning abilities ouyang2022training ; ziegler2019fine ; rafailov2023direct ; guo2024direct ; shao2024deepseekmath . A family of successful algorithms, including PPO ziegler2019fine , DPO rafailov2023direct , online DPO guo2024direct , and GRPO shao2024deepseekmath , is now being actively extended to LMMs, although applications have mainly focused on image-LMMs sun2023aligning ; yu2024rlhf ; zhang2025r1 ; zhang2024improve , with video comprehension being a newer frontier. While these efforts have concentrated on mitigating visual hallucination zhang2024direct ; wang2024videohallucer ; zhang2024eventhallusion or (concurrently) enabling structured reasoning feng2025videor1 ; chen2025exploring ; luo2025thinking like Chain-of-Thought wei2022chain , our work investigates a distinct dimension: enhancing the model’s fundamental temporal awareness, by leveraging AoT signals naturally embedded in videos.
@@ 3.1 Are Today’s Benchmarks Time-Sensitive? Revisiting Current LMM Evaluation
We then evaluate the three selected LMMs across eight popular VQA benchmarks (VITATECS li2024vitatecs , TemporalBench cai2024temporalbench , NExT-QA xiao2021nextqa , PerceptionTest patraucean2023perception , VideoMME fu2024videomme , Vinoground zhang2024vinoground , TempCompass liu2024tempcompass and TVBench cores2024tvbench ). [^3] Note that our analysis focuses on MCQ tasks for their popularity and clear evaluation structure, as opposed to open-ended QA or captioning tasks that require an additional LLM evaluator, introducing uncertainty. To quantify benchmark and sample-level AoT sensitivity (facilitating later construction of AoTBench, c.f., Sec. 3.3), we propose the temporal divergence score (TDS). For an evaluator LMM $f$, consider an MCQ sample $i$, comprised of video $v_{i}$, language query $l_{i}$ and ground truth answer $y_{i}$ from $K_{i}$ options. Denote the reversed video as $\tilde{v}_{i}$. The MCQ setup instructs the model to respond only with the option’s letter. The first-token probability distribution is therefore a direct measure of the model’s confidence over the $K_{i}$ choices brown2020language ; hendrycks2020measuring ; lin2021truthfulqa . We compute the probability distribution for both forward and reversed video: $p_{i}=\mathrm{FirstTokenProb}(f,v_{i},l_{i})$ and $\tilde{p}_{i}=\mathrm{FirstTokenProb}(f,\tilde{v}_{i},l_{i})$, where $p_{i},\tilde{p}_{i}\in\mathbb{R}^{K_{i}}$.To ensure our analysis focuses on instances genuinely requiring arrow of time (AoT) reasoning, we first filter out outlier samples where performance on the reversed video exceeds that of the forward one, as these cases typically indicate ambiguous or irrelevant temporal cues. The TDS for any remaining sample $i$ is defined using KL divergence: $\mathrm{TDS}_{i}=\mathbb{D}_{\text{KL}}[p_{i}\|\tilde{p}_{i}]$. A high TDS signifies that the model’s prediction is highly sensitive to the video’s temporal direction—causing disagreement between forward and reverse responses—therefore suggesting the sample requires AoT understanding for a correct interpretation and answer.
@@ 3.2 ArrowRL: Learning the Arrow of Time
We now discuss our approach to equipping a given LMM $\pi_{\theta}$ with AoT understanding for enhanced temporal perception. Our core premise is that videos are inherently dynamic, and LMMs should capture this dynamic progression when responding. Specifically, given a question $q=(v,l)$ (composed of video $v$ and language query $l$) and a target response $o^{\star}$, drawn from the underlying data distribution $P(Q,O^{\star})$ of a standard instruction tuning dataset, the goal is for $\pi_{\theta}$ to produce responses that align well with $o^{\star}$. To foster broad AoT sensitivity, our training utilizes a diverse suite of tasks—MCQ, open-ended QA, and video captioning (see Sec. 4 for the specific datasets used).
@@
First, we define a fidelity reward encouraging alignment between each candidate response $o_{i}$ and the target response $o^{\star}$, $r_{i}^{\mathrm{\ fid}}=\mathrm{Similarity}(o_{i},o^{\star})$. The function $\mathrm{Similarity}(\cdot,\cdot)$ returns a scalar between 0 and 1 and is calculated via accuracy matching (1.0 if $o_{i}=o^{\star}$, else 0.0) for MCQ tasks, or via an LLM-based similarity score for open-ended QA and captioning (see Supp. B.3 for the specific prompts and the LLM judge employed).
@@
[…]
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\$$\mathbb{E}_{(q,o^{\star})\sim P(Q,O^{\star}),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg\{\min\Bigg[\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}\hat{A}_{i,t},$
@@ 4 Experiments
Our ArrowRL training data comprises a comprehensive suite of tasks, with data selected or adapted for each to emphasize scenarios requiring AoT awareness: (1) MCQ tasks: we format sequence direction classification (Fig. 1 (a)) as MCQ, using 1.1K training videos selected from UCF101 soomro2012ucf101 (following prior use of this dataset ghodrati2018videotime ); (2) Open-ended QA: we curate a high-temporality subset of LLaVA-Video-178K zhang2024llavavideo , filtering based on the perplexity difference between forward and reverse video to retain samples with great temporal sensitivity, totaling 11.8K samples; (3) Video Captioning: we employ the training set du2024reversed of RTime, which provides high-temporality videos alongside distinct human captions for their forward and reverse versions, comprising 11.7K samples. See Supp. B.3 for data examples and prompts used.
@@
Our framework presents a straightforward and highly efficient approach by enabling direct application to pretrained LMMs without an intermediate supervised fine-tuning (SFT) step; we find ArrowRL alone is sufficient to instill AoT awareness. We demonstrate its broad applicability across three leading base LMMs: LLaVA-OV-7B li2024llavaov , Qwen2-VL-7B wang2024qwen2 , and Qwen2.5-VL-7B bai2025qwen25 . For efficiency and memory, we limit the maximum number of video frames to 16 during training. Hyperparameter $\alpha$ is set as 0.25 and $\gamma$ is set as 0.75. The response group size $G$ is set as 8. Training consists of 2000 RL steps on 6 NVIDIA GH200 GPUs.
@@
Main Results
As shown in Table 1, ArrowRL effectively enhances models’ temporal perception capabilities across diverse settings. Results on AoTBench (left table) reveal critical deficiencies in current LMM AoT perception. Visually-focused tasks like sequence direction classification prove especially challenging: nearly all open-source models exhibit a universal failure mode: their responses are identical regardless of video input, resulting in consistent chance-level performance (50.0% on ReverseFilm and 51.6% on UCF101). Leading proprietary models (GPT-4o, Gemini-1.5-Pro), while slightly better, still perform far below satisfactory levels. This is in stark contrast to the proprietary models’ absolute accuracy on other mainstream VQA tasks, where accuracy hovers around 70-85%. Our finding and the proposed AoTBench illuminate visual temporal bottlenecks. Importantly, this improved AoT awareness also translates to gains on existing, unfiltered temporal VQA benchmarks (right table), such as +10.8 group score (+65.9% relative gain) for Qwen2.5-VL-7B on Vinoground. These results underscore the promise of enhancing AoT perception for overall temporal perception. See Supp. C.1 for a follow-up discussion of the relative strengths of different base models and how they translate to ArrowRL’s largest margins.
@@
Table 2: Results on general video benchmarks (base vs. Arrow-RL enhanced LLaVA-OV-7B). ArrowRL’s specialization on temporal awareness does not degrade performance on these general video understanding tasks.
Model VideoMME (S) NExT-QA TemporalBench (S) VITATECS PerceptionTest
LLaVA-OV-7B li2024llavaov 67.78 77.11 60.57 84.66 57.48
+ ArrowRL 68.11 78.11 61.17 85.37 57.64
Results on General Video Benchmarks
To ensure that specializing in AoT does not degrade general video understanding, we evaluate ArrowRL-enhanced LLaVA-OV-7B on five benchmarks previously identified as less temporally sensitive (see Fig. 3 right). As shown in Table 2, ArrowRL preserves or even slightly improves performance on these general benchmarks where AoT is not required, indicating it enhances temporal perception without sacrificing general video understanding.
Table 3: Ablation results of ArrowRL. We report average performance across all columns of AoTBench tasks in Table 1. ArrowRL greatly outperforms the SFT baseline trained on the same data, demonstrating the effectiveness of our RL approach. In addition, using our curated high-temporality post-training data provides a performance gain, validating our data selection strategy.
Model Training data Acc. (%)
Qwen2.5-VL-7B bai2025qwen25 - 56.2
+ SFT ArrowRL post-train set 57.4
+ ArrowRL LLaVA-Video-178K captions 57.7
+ ArrowRL RTime captions 60.4
+ ArrowRL ArrowRL post-train set 61.4
Figure 7: Hyperparameter analysis of ArrowRL on AoTBench, using Qwen2.5-VL-7B as the base LMM. We identify an optimal setting at $\alpha=0.25$ (reverse reward weight, left) and $\gamma=0.75$ (dynamic weighting threshold, right), which effectively balances the fidelity and reverse reward.
Ablations
Fig. 7 presents our hyperparameter analysis, confirming two key design choices. First, the reverse reward is essential: the left plot shows that removing it ($\alpha=0$) degrades performance to below that of the base LMM. Second, the dynamic weighting mechanism is beneficial: the right plot shows that disabling it ($\gamma=1$) yields less improvement than our dynamic approach, which can zero out the reward for non-AoT-sensitive samples. We further investigate different inference frame settings and find ArrowRL generalizes robustly despite its fixed-frame training. See Supp. C.1 for details.
@@
Fig. 8 provides qualitative examples, comparing base LMMs against our ArrowRL enhancement. The VQA examples (left) demonstrate ArrowRL enabling base models to correctly answer temporally challenging questions. Similarly, when prompted to describe forward and reverse sequences (right), the base model provides the “unwrapping” description for both, failing to connect frames dynamically to understand the temporal progression and possibly influenced by language bias (favoring the more common “unwrapping”). Conversely, ArrowRL demonstrates clear temporal awareness by producing distinct and correct captions reflecting the temporal direction. See Supp. C.2 for more qualitatives, including failure cases.
Figure 8: Qualitative results comparing base LMMs vs. Arrow-RL enhanced models. ArrowRL correctly answers AoT-sensitive VQA questions (left) and generates distinct, temporally accurate captions (right), while the base model fails to capture temporal progression in these examples. More examples and failure cases in Supp. C.2.
Limitations and Future Work
While AoTBench grounds temporal directionality in language, we acknowledge that describing irreversible events remains inherently difficult in natural language. Future work could explore how learned representations encode time beyond language bagad2025chirality , and probe whether models develop specialized activations for time-reversed, physically implausible inputs deco2023arrow . For ArrowRL, important future directions include extending our methods to long videos (e.g., via keyframe selection) and exploring integration with temporal reasoning tasks that necessitate Chain-of-Thought reasoning. See Supp. D for a comprehensive discussion of limitations.